What is the quantity and how to dissolve silver chromate in nitric acid

1
Expert's answer
2022-05-11T07:56:03-0400

The chemical reaction of the silver chromate dissolution is:


Ag2CrO4 + 2HNO3 → 2AgNO3 + H2O + CrO3


The silver chromate is sensitive to pH. In water the pH is neutral, so the crystals are not dissolving, in a basic pH there are precipitates of silver oxide and hydroxide. In acidic pH it is soluble in dilute nitric acid because the products of this reaction, silver nitrate (AgNO3) and chromic acid CrOare both soluble.


The molar solubility of silver chromate depends on the nitric acid concentration, so the desired answer cannot be precise. The "K_{sp} = 1.1 \times 10^{-12}"

"Ag_2CrO_4 = 2Ag^+ + CrO_4^{2-}"

At equilibrium the "K_{sp}" is:

"1.1 \times 10^{-12} = x\times (2x)^2 = 4 x^3"

"x = 6.5 \times 10^{-5} M"

Where x is the concentration of silver chromate. Knowing the volume of nitric acid, the quantity of silver chromate may be deduced.